# Fine-tuning an Large Language Model for Automating Computational Fluid Dynamics Simulations

### v1

@@ Abstract
Configuring computational fluid dynamics (CFD) simulations typically demands extensive domain expertise, limiting broader access. Although large language models (LLMs) have advanced scientific computing, their use in automating CFD workflows is underdeveloped. We introduce a novel approach centered on domain-specific LLM adaptation. By fine-tuning Qwen2.5-7B-Instruct on NL2FOAM, our custom dataset of 28716 natural language-to-OpenFOAM configuration pairs with chain-of-thought (CoT) annotations, we enable direct translation from natural language descriptions to executable CFD setups. A multi-agent framework orchestrates the process, autonomously verifying inputs, generating configurations, running simulations, and correcting errors. Evaluation on a benchmark of 21 diverse flow cases demonstrates state-of-the-art performance, achieving 88.7% solution accuracy and 82.6% first-attempt success rate. This significantly outperforms larger general-purpose models like Qwen2.5-72B-Instruct, DeepSeek-R1, and Llama3.3-70B-Instruct, while also requiring fewer correction iterations and maintaining high computational efficiency. The results highlight the critical role of domain-specific adaptation in deploying LLM assistants for complex engineering workflows.
@@ 1 Introduction
LLMs demonstrate remarkable natural language capabilities across diverse domains [8, 9, 10, 11]. They have aided scientific discovery [12] in fields like mathematics [13] and chemistry [14]. However, their application to computational physics, particularly CFD, is constrained by the need for precise physical understanding and software-specific syntax. While general-purpose LLMs like GPT-4 [8] and DeepSeek [10] contain broad scientific knowledge, they lack the specialized expertise crucial for reliable CFD automation, often generating physically inconsistent parameters or syntactically incorrect configurations. This domain knowledge gap hinders effective automation of the complex CFD workflow, which demands deep understanding of numerical schemes, turbulence models, boundary conditions, and solver-specific implementation details. Addressing this challenge requires domain-specific adaptation that aligns LLM capabilities with the rigorous requirements of CFD.
@@
Retrieval-augmented generation (RAG) [15] has been proposed to address the domain knowledge gap in CFD automation by allowing general-purpose LLMs to access specialized information. Examples include MetaOpenFOAM [16], a RAG-based multi-agent system that automates OpenFOAM simulation workflows from natural language inputs and later extended to post-processing [17]. Similarly, Pandey et al[18] developed RAG-based OpenFOAMGPT, demonstrating zero-shot case setup and condition modification capabilities with GPT-4o and GPT-o1. The OpenFOAMGPT framework was then utilized to evaluate the cost-effectiveness of different LLMs [19]. However, RAG-based systems are inherently limited by their reliance on knowledge retrieval; they access pre-existing information rather than developing true domain understanding [20, 21]. This limitation becomes particularly problematic for new configurations or complex physical scenarios absent from reference materials, potentially leading to in fragmented knowledge integration and potentially physically inconsistent outputs [22, 23]. Furthermore, selecting optimal numerical configurations in CFD is challenging as settings are often non-unique and highly dependent on specific flow regimes and geometric complexities—a nuance difficult to capture through simple retrieval. These shortcomings underscore the need for more sophisticated approaches that embed domain expertise directly into model parameters through specialized fine-tuning.
@@
Fine-tuning [24] offers a direct approach for embedding domain expertise into LLMs, potentially overcoming RAG’s limitations by incorporating CFD knowledge into the model’s parameters rather than relying on external retrieval. It enables models to develop deeper understanding of fluid dynamics principles, numerical methods, and solver-specific requirements. Effective fine-tuning requires high-quality labeled datasets [25] specifically tailored to CFD applications—a challenge given the complexity and diversity of simulation scenarios. Such datasets must capture not only correct syntax and parameter settings but also the underlying physical reasoning and problem-specific considerations that guide expert decisions. Despite the challenges, fine-tuned models offer potentially greater consistency in handling new cases, enhanced physics reasoning, and robust performance without the computational overhead and latency associated with retrieval systems [26].
@@
In this work, we developed a domain-specific fine-tuned LLM for automating CFD workflows. We developed NL2FOAM, a custom dataset comprising 28716 pairs of natural language descriptions and corresponding OpenFOAM configurations, augmented with chain-of-thought (CoT) annotations to capture expert reasoning. Using NL2FOAM, we fine-tuned Qwen2.5-7B-Instruct [9], enabling it to translate high-level natural language problem descriptions into executable CFD setups. A multi-agent framework manages the workflow, handling input verification, configuration generation, simulation execution, and error correction autonomously. Our evaluation on a benchmark of 21 diverse flow cases demonstrates state-of-the-art performance, significantly surpassing larger general-purpose models and highlighting the effectiveness of specialized fine-tuning for complex engineering tasks.
@@ 2.1 Fine-tuning LLM
LLMs, pre-trained on vast corpora of text, encode broad knowledge and language capabilities [8, 10, 9, 11]. They can be specialized for domain-specific applications through fine-tuning [25], particularly when sufficient labeled training data exists and high-precision is essential. Conventional fine-tuning updates all model parameters, imposing significant computational burden. Low-rank adaptation (LoRA) [27] substantially reduces the computational footprint while maintaining comparable performance. Mathematically, LoRA updates a pre-trained weight matrix $W\in\mathbb{R}^{d\times k}$ by adding $\delta W=BA$ where $B\in\mathbb{R}^{d\times r}$, $A\in\mathbb{R}^{r\times k}$, and the rank $r\ll\min(d,k)$. This approach typically reduces tunable parameters by 100- to 1000-fold. During inference, the updated weight $W^{\prime}=W+\delta W$ is used efficiently without additional computational overhead.
For the CFD-specific adaptation, we fine-tuned Qwen2.5-7B-Instruct [9] via LoRA on the NL2FOAM dataset, which comprises 28716 natural language-to-OpenFOAM configuration cases. Each case maps input, a natural language description, mesh files, and input file templates, to outputs including numerical configurations, initial fields, boundary conditions, and an execution script. The output is annotated with CoT [28] reasoning to enhance generation quality and stability, which is validated through an ablation study. Details of NL2FOAM will be introduced in Sec. 2.3. In practical applications, input file templates are generated automatically based on natural language descriptions and mesh files. The resulting fine-tuned LLM functions as an intelligent interface between users and OpenFOAM, translating natural language specifications into executable CFD configurations without requiring users to master OpenFOAM’s complex syntax and parameter structures.
We applied LoRA with a rank $r=8$, reducing trainable parameters from 7.6B to 0.02B. Fine-tuning utilized Llamma-Factory [29], a unified framework enabling efficient and flexible LLM adaptation. We trained the model using four NVIDIA GeForce RTX 4090 GPUs, employing AdamW [30] optimization. The baseline learning rate was $5\times 10^{-5}$, with linear warmup over the first 10% of steps to mitigate initial instability. A total batch size of 16 balanced GPU memory constraints and training efficiency. Training spanned four epoches, checkpointing at each epoch boundary; this duration was empirically determined from validation loss plateaus to balance convergence and computational cost.
@@ 2.2 Multi-agent system
The fine-tuned LLM serves as the core reasoning engine in our multi-agent framework. This framework orchestrates CFD workflow automation and enhances domain-specific reasoning, building upon previous RAG approaches [16]. As illustrated in Fig. 1, the system orchestrates four specialized agents—pre-checker, LLM generator, runner, and corrector—through a structured workflow that enforces OpenFOAM syntax compliance and numerical stability.
Figure 1: The multi-agent framework for automated CFD simulation, centered on a fine-tuned LLM. The workflow progresses from user input through a pre-checker, LLM-based generation of input files, simulation execution by the runner, and an iterative error correction loop involving the corrector, ultimately producing simulation results.
The workflow begins with a user’s natural language description of the CFD problem and mesh files. The pre-checker validates inputs, queries users if needed, and generates input templates incorporating boundary names extracted from the mesh files. The fine-tuned LLM then generates the OpenFOAM case directory through structured CoT reasoning, including numerical configurations, initial fields, boundary conditions, and an execution script. The runner executes simulations while monitoring real-time logs. If errors occur, the corrector analyzes and resolves issues. The corrected files are then resubmitted to the runner, continuing this cycle until the simulation completes successfully.
We employ Qwen2.5-72B-Instruct [9] for the pre-checker, runner, and corrector agents. The multi-agent architecture utilizes MetaGPT v0.8.1 [31], leveraging its metaprogramming capabilities for rapid system prototyping. Simulations run on OpenFOAM-v2406 [7]. By integrating linguistic understanding with numerical rigor, this approach reduces manual configuration efforts while maintaining engineering validity.
@@ 2.3 NL2FOAM
As sketched in Fig. 2, we built the NL2FOAM dataset through a semi-automated pipeline to fine-tune an LLM for CFD automation. The dataset contains 28716 cases that link natural language descriptions with executable OpenFOAM configurations, each including a CFD problem description, mesh files, input files (numerical configurations, initial fields, and boundary conditions), an execution script, and a structured CoT reasoning trace.
@@
Starting with 16 OpenFOAM cases (see Tab. 1) spanning laminar and turbulent flows with the Reynolds number ($Re$) from $10$ to $9\times 10^{7}$, we modified input files (controlDict, fvScheme, and fvSolution) to generate over 100k variations. An LLM enhanced linguistic diversity by rephrasing the problem descriptions. Subsequently, comprehensive simulation testing filtered out approximately 72% of cases with runtime errors, solution divergence, or excessive runtime (more than 72 hours). CoT reasoning traces were then generated for the valid cases. This construction process balances physical accuracy with linguistic diversity by combining expert-curated foundations and systematic augmentation.
@@
We standardized LLM inputs using structured prompts including a natural language CFD problem description, mesh files, and json input templates. The description details flow configuration, boundary conditions, and solver parameters (e.g., “Conduct a RANS simulation using k-epsilon model for a cylinder wake with inlet velocity of 10 m/s”). The mesh files ensures geometric consistency and guide boundary condition assignments. The json templates define required input files (e.g., 0/U, system/controlDict) with placeholders for boundary conditions derived from mesh files. These templates are automatically generated through the multi-agent system, do not require user input in practice.
Our fine-tuning utilized synthetic CoT solutions [28], known to enhanced small model performance [32]. Each prompt is paired with a ground truth solution featuring a CoT annotation with six reasoning steps: (1) define the problem and objectives; (2) select the appropriate solver and turbulence models; determine the required OpenFOAM files; (4) apply the boundary and initial conditions; (5) configure the simulation parameters; (6) write the execution script.
The OpenFOAM configuration files and the execution script follow the CoT to form the full ground truth. An example case from NL2FOAM is shown in Fig. 3.
@@
# $<$input templates$>$
# Response: # $<$CoT reasoning steps$>$ # Based on the above thought process, the complete Foam files and Allrun scripts are as follows: ## $<$OpenFOAM files$>$
## $<$execution script$>$
Figure 3: Structure of an example case within the NL2FOAM dataset used for fine-tuning. The input consists of a natural language description (Question) and placeholders for mesh files and templates. The output (Response) includes CoT reasoning steps followed by the generated OpenFOAM files and execution script.
@@ 2.4 Benchmark
We evaluate our framework using an incompressible flow benchmark of 21 cases (10 laminar, 11 turbulent) with $Re$ spanning from $40$ to $5\times 10^{8}$, as listed in Tab. 2. 71% of test cases (15/21) extend beyond the OpenFOAM tutorial, including a jet flow and turbulent square column wakes. To ensure a fair assessment, there is no configuration overlap between the training and benchmark sets. Although two case names appear in both sets (cylinder wake and Poiseuille flow), their parameters differ substantially. While the training set include a laminar cylinder wake at $Re=40$, the benchmark tests this geometry from $Re=40$ to $1\times 10^{5}$ (excluding 100). Likewise, the Poiseuille flow parameters change from $Re=1500$ (training) to $Re=100$ (benchmark). This separation ensures the evaluation assesses generalization across diverse $Re$ and flow regimes, not memorization. Furthermore, the benchmark includes multi-solver configurations (e.g., cylinder wake validated with icoFoam, simpleFoam, pisoFoam, and pimpleFoam) to test the framework’s ability to select context-appropriate numerical methods.
@@
We evaluated performance using five metrics: accuracy, pass@1, iterations, token usage, and expense. An “experiment” proceeds from inputting a natural language description and mesh files to obtaining CFD simulation results. An experiment “passes” if it achieves a convergent solution in 72 hours at most 10 correction attempts; otherwise, it “fails”. Each benchmark case undergoes $n=10$ independent experiments. The final reported metrics are averaged across all experiments, which are listed below.
[…]

### v2

# Fine-tuning a Large Language Model for Automating Computational Fluid Dynamics Simulations

@@ Abstract
Configuring computational fluid dynamics (CFD) simulations typically demands extensive domain expertise, limiting broader access. Although large language models (LLMs) have advanced scientific computing, their use in automating CFD workflows is underdeveloped. We introduce a novel approach centered on domain-specific LLM adaptation. By fine-tuning Qwen2.5-7B-Instruct on NL2FOAM, our custom dataset of 28716 natural language-to-OpenFOAM configuration pairs with chain-of-thought (CoT) annotations enables direct translation from natural language descriptions to executable CFD setups. A multi-agent system orchestrates the process, autonomously verifying inputs, generating configurations, running simulations, and correcting errors. Evaluation on a benchmark of 21 diverse flow cases demonstrates state-of-the-art performance, achieving 88.7% solution accuracy and 82.6% first-attempt success rate. This significantly outperforms larger general-purpose models like Qwen2.5-72B-Instruct, DeepSeek-R1, and Llama3.3-70B-Instruct, while also requiring fewer correction iterations and maintaining high computational efficiency. The results highlight the critical role of domain-specific adaptation in deploying LLM assistants for complex engineering workflows. Our code and fine-tuned model have been deposited at https://github.com/YYgroup/AutoCFD.
@@ 1 Introduction
LLMs demonstrate remarkable natural language capabilities across diverse domains [8, 9, 10, 11]. They have aided scientific discovery [12] in fields like mathematics [13] and chemistry [14]. However, their application to computational physics, particularly CFD, is constrained by the need for precise physical understanding and software-specific syntax. While general-purpose LLMs like GPT-4 [8] and DeepSeek-R1 [10] contain broad scientific knowledge, they lack the specialized expertise crucial for reliable CFD automation, often generating physically inconsistent parameters or syntactically incorrect configurations. This domain knowledge gap hinders effective automation of the complex CFD workflow, which demands a deep understanding of numerical schemes, turbulence models, boundary conditions, and solver-specific implementation details. Addressing this challenge requires domain-specific adaptation that aligns LLM capabilities with the rigorous requirements of CFD.
@@
Retrieval-augmented generation (RAG) [15] has been proposed to address the domain knowledge gap in CFD automation by allowing general-purpose LLMs to access specialized information. Examples include MetaOpenFOAM [16], a RAG-based multi-agent system that automates OpenFOAM simulation workflows from natural language inputs and later extended to post-processing [17]. Similarly, Pandey et al[18] developed RAG-based OpenFOAMGPT, demonstrating zero-shot case setup and condition modification capabilities with GPT-4o and GPT-o1. The OpenFOAMGPT framework was then utilized to evaluate the cost-effectiveness of different LLMs [19]. However, RAG-based systems are inherently limited by their reliance on knowledge retrieval; they access pre-existing information rather than developing true domain understanding [20, 21]. This limitation becomes particularly problematic for new configurations or complex physical scenarios absent from reference materials, potentially leading to fragmented knowledge integration and physically inconsistent outputs [22, 23]. Furthermore, selecting optimal numerical configurations in CFD is challenging as settings are often non-unique and highly dependent on specific flow regimes and geometric complexities—a nuance difficult to capture through simple retrieval. These shortcomings underscore the need to embed domain expertise directly into model parameters.
@@
Fine-tuning [24] offers a direct approach for embedding domain expertise into LLMs, potentially overcoming RAG’s limitations by incorporating CFD knowledge into the model’s parameters rather than relying on external retrieval. It enables models to develop a deeper understanding of fluid dynamics principles, numerical methods, and solver-specific requirements. Effective fine-tuning requires high-quality labeled datasets [25] specifically tailored to CFD applications—a challenge given the complexity and diversity of simulation scenarios. Such datasets must capture not only correct syntax and parameter settings but also the underlying physical reasoning and problem-specific considerations that guide expert decisions. Despite the challenges, fine-tuned models offer potentially greater consistency in handling new cases, enhanced physics reasoning, and robust performance without the computational overhead and latency associated with retrieval systems [26].
@@
In this work, we developed a domain-specific fine-tuned LLM for automating CFD workflows. We developed NL2FOAM, a custom dataset comprising 28716 pairs of natural language descriptions and corresponding OpenFOAM configurations, augmented with chain-of-thought (CoT) annotations to capture expert reasoning. Using NL2FOAM, we fine-tuned Qwen2.5-7B-Instruct [9], enabling it to translate high-level natural language problem descriptions into executable CFD setups. A multi-agent system manages the workflow, handling input verification, configuration generation, simulation execution, and error correction autonomously. Our evaluation on a benchmark of 21 diverse flow cases demonstrates state-of-the-art performance, significantly surpassing larger general-purpose models and highlighting the effectiveness of specialized fine-tuning for complex engineering tasks.
@@ 2.1 Fine-tuning LLM
LLMs, pre-trained on vast corpora of text, encode broad knowledge and language capabilities [8, 10, 9, 11]. They can be specialized for domain-specific applications through fine-tuning [25], particularly when sufficient labeled training data exists and high-precision is essential. Conventional fine-tuning updates all model parameters, imposing significant computational burden. Low-rank adaptation (LoRA) [27] substantially reduces the computational footprint while maintaining comparable performance. Mathematically, LoRA updates a pre-trained weight matrix $W\in\mathbb{R}^{d\times k}$ by adding $\delta W=BA$ with $B\in\mathbb{R}^{d\times r}$, $A\in\mathbb{R}^{r\times k}$, and the rank $r\ll\min(d,k)$. This approach typically reduces tunable parameters by 100- to 1000-fold. During inference, the updated weight $W^{\prime}=W+\delta W$ is used efficiently without additional computational overhead.
For the CFD-specific adaptation, we fine-tuned Qwen2.5-7B-Instruct [9] via LoRA on the NL2FOAM dataset, which comprises 28716 natural language-to-OpenFOAM configuration cases. Details of NL2FOAM will be introduced in Sec. 2.3, and fine-tuning specifics are provided in A. In practice, the model processes a natural language description (e.g., “Conduct a RANS simulation using k-epsilon model for a cylinder wake with inlet velocity of 10 m/s”), mesh files, and input file templates. The mesh files are required as LLMs cannot reliably generate them, while templates are generated automatically from the description and mesh files. Although the fine-tuned model generally infers appropriate initial flow fields, explicit specification in the description improves success rates. The resulting fine-tuned LLM functions as an intelligent interface, translating user descriptions into executable OpenFOAM configurations without requiring expertise in its syntax and parameter structures.
@@ 2.2 Multi-agent system
The fine-tuned LLM serves as the core reasoning engine in our multi-agent system. This system orchestrates CFD workflow automation and enhances domain-specific reasoning, building upon previous RAG approaches [16]. As illustrated in Fig. 1, this system orchestrates four specialized agents—pre-checker, LLM generator, runner, and corrector—through a structured workflow that enforces OpenFOAM syntax compliance and numerical stability. Further details on the multi-agent system setup can be found in the supplementary material.
Figure 1: The multi-agent system for automated CFD simulation, centered on a fine-tuned LLM. The workflow progresses from user input through a pre-checker, LLM-based generation of input files, simulation execution by the runner, and an iterative error correction loop involving the corrector, ultimately producing simulation results.
The multi-agent workflow begins with a user’s natural language description of the CFD problem and mesh files. The pre-checker validates inputs, queries users if needed, and generates input templates incorporating boundary names extracted from the mesh files. The fine-tuned LLM then generates the OpenFOAM case directory through structured CoT reasoning, including numerical configurations, initial fields, boundary conditions, and an execution script. The runner executes simulations while monitoring real-time logs. If errors occur, the corrector analyzes and resolves issues. The corrected files are then resubmitted to the runner, continuing this cycle until the simulation completes successfully.
@@ 2.3 NL2FOAM
As sketched in Fig. 2, we built the NL2FOAM dataset to fine-tune an LLM for CFD automation. Starting with 16 OpenFOAM cases (see Tab. 1) spanning laminar and turbulent flows with the Reynolds number ($Re$) from $10$ to $9\times 10^{7}$, we modified configuration files to generate over 100k variations. An LLM enhanced linguistic diversity by rephrasing the problem descriptions. Test simulations filtered out cases with runtime errors, solution divergence, or excessive runtime (more than 72 hours). The final dataset contains 28716 cases that link natural language descriptions to executable OpenFOAM configurations, each including a problem description, mesh files, OpenFOAM input files (numerical configurations, initial fields, and boundary conditions), an execution script, and an LLM-generated CoT reasoning trace.
@@
The NL2FOAM dataset exhibits diversity across multiple dimensions. We selected various OpenFOAM cases covering steady and unsteady flows, laminar and turbulent regimes, and multiple simulation methods (directly solving the Navier-Stokes equations for laminar flows, and Reynolds-averaged Navier-Stokes simulations and large-eddy simulations for turbulent flows). We further expanded diversity through data augmentation, varying numerical parameters including finite volume schemes, velocity-pressure coupling solvers, solver parameters, and linear system solvers. Additionally, LLM-based rephrasing of problem descriptions ensures linguistic diversity in the natural language specifications. Quality is assured through test simulations, eliminating invalid configurations from the final dataset.
Despite these strengths, the dataset is limited to incompressible flow cases, excluding more complex phenomena such as multiphase, compressible, or reacting flows, which often require additional fields and distinct solver settings. Additionally, the limited number of base OpenFOAM cases restricts coverage even within incompressible flows. In the present work, we limited our initial scope to incompressible flows to establish a robust foundation. Future work should systematically expand to include more complex flows, incorporating domain-specific constraints.
Each training case in NL2FOAM is a string pairing a structured input with a ground truth response, as illustrated in Fig. 3. The input includes a natural language problem description, mesh file references, and input templates. The ground truth contains complete OpenFOAM configurations and an execution script, preceded by CoT reasoning steps. We utilized the synthetic CoT solutions [28] to enhance small model performance [29]. These ground truth CoT annotations are generated by a separate LLM, following a six-step reasoning structure provided in the prompt: (1) define the problem and objectives; (2) select the appropriate solver and turbulence models; (3) determine the required OpenFOAM files; (4) apply the boundary and initial conditions; (5) configure the simulation parameters; (6) write the execution script. The final OpenFOAM files and script adhere to this reasoning.
@@
# Input file template:
    "0/p": {
        "boundaryField": {
            "inlet": { "type": "xxx", ... },
# Response: # CoT reasoning steps: ## 1. define the problem and objectives: … ## 2. select the appropriate solver and turbulence models: … ## 3. determine the required OpenFOAM files: … ## 4. apply the boundary and initial conditions: … ## 5. configure the simulation parameters: … ## 6. write the execution script: … # Based on the above thought process, the complete foam files and execution scripts are as follows: ## OpenFOAM files
## execution script
Figure 3: Structure of an example case within the NL2FOAM dataset used for fine-tuning. The input consists of a natural language description (Question), mesh file references, and input templates. The output (Response) includes CoT reasoning steps followed by the generated OpenFOAM files and execution script.
@@ 2.4 Benchmark
We evaluate our framework using an incompressible flow benchmark of 10 laminar and 11 turbulent cases with $Re$ spanning from $40$ to $5\times 10^{8}$, as listed in Tab. 2. 71% of test cases (15/21) extend beyond the OpenFOAM tutorial, including a jet flow and turbulent square column wakes. To ensure a fair assessment, there is no configuration overlap between the training and benchmark sets. Although two case names appear in both sets (cylinder wake and Poiseuille flow), their parameters differ substantially. While the training set includes a laminar cylinder wake at $Re=100$, the benchmark tests this geometry from $Re=40$ to $1\times 10^{5}$ (excluding 100). Likewise, the Poiseuille flow parameters change from $Re=1500$ (training) to $Re=100$ (benchmark). This separation ensures the evaluation assesses generalization across diverse $Re$ and flow regimes, not memorization. Furthermore, the benchmark includes multi-solver configurations (e.g., cylinder wake validated with icoFoam, simpleFoam, pisoFoam, and pimpleFoam) to test the framework’s ability to select context-appropriate numerical methods.
@@
We evaluated performance using five metrics: accuracy, pass@1, iterations, token usage, and expense. An “experiment” proceeds from inputting a natural language description and mesh files to obtaining CFD simulation results. An experiment “passes” if it achieves a convergent solution in 72 hours at most 10 correction attempts; otherwise, it “fails”. Each benchmark case undergoes $n=10$ independent experiments. The final reported metrics are averaged across all experiments.
“Accuracy” measures solution reliability using the L2 norm $\epsilon$ between the LLM-based automated CFD solution and the benchmark, defined as $1-\epsilon$. Failed experiments receive 0% accuracy. “pass@$1$” [30] represents the first-attempt success rate, defined as the ratio of the number of successful experiments to the total number of attempts (10 attempts in this study). “Iterations” count the correction rounds needed to fulfill user requirements, indicating convergence efficiency. “Token Usage” measures LLM tokens consumed, reflecting computational resource requirements. “Expense” quantifies actual costs per experiment, differing from token usage due to varying prices.